# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v1

@@ Appendix C Student Objective
Our aim is to train a masked-diffusion student whose forward-time marginals match those of a teacher across the entire time horizon. We adopt the Integral KL (IKL) between the student and teacher forward marginals $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ as the training objective. In the masked setting, the forward-time corruption kernel is a forward absorbing process, e.g.,
@@
$\nabla_{\nu}\mathcal{L}(\nu)\approx\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\omega\left(t_{i}\right)}{\pi\left(t_{i}\right)}\cdot R\left(\mathbf{z}_{t_{i}},t_{i}\right)\cdot\nabla_{\nu}\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]$ (15)
@@ Appendix D Auxiliary Discriminator for Density Ratio Estimation
Our student objective, as derived in Appendix C, relies on the reward term $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$, which is intractable due to the unknown marginal distributions $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. To overcome this, we introduce a tractable estimator for the density ratio $\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t})}{\mathbf{q}_{\theta}(\mathbf{z}_{t})}$ by training an auxiliary discriminator network. This approach is inspired by the principles of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Wang et al., 2023) and has been successfully applied in (Wang et al., 2025). We hereby adapt the following lemma to MDM setting.
For clarity, we adopt a time-agnostic notation where the corruption level is parameterized by the masking ratio $t\in[0,1]$, which corresponds to the schedule $\alpha_{t}$ via $t=1-\alpha_{t}$. Let $\mathbf{z}_{t}\in\mathcal{V}^{L}$ denote a sequence with a mask ratio at time $t$.
Lemma D.1 (Density Ratio Representation).
Let $\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ and $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ be the teacher and student models respectively, over the discrete state space $\mathcal{V}^{L}$ at a mask ratio $t$. Consider a discriminator $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$, which outputs a probability for each position $\ell\in\{1,\dots,L\}$. The discriminator $D_{\lambda}$ is trained to minimize the objective
@@
We here provide a detailed derivation for the optimal discriminator in the context of our discrete state space. The objective $\mathcal{J}(D)$ is an expectation over the discrete random variable $\mathbf{z}_{t}$. We can express this expectation as a summation over all possible sequences $\mathbf{z}_{t}\in\mathcal{V}^{L}$:
$\mathcal{J}(D)=\sum_{\mathbf{z}_{t}\in\mathcal{V}^{L}}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\log D_{\lambda}(\mathbf{z}_{t},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\log(1-D_{\lambda}(\mathbf{z}_{t},t))\right]$
@@
$\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}$ $\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})\log D_{\lambda}(\mathbf{z}_{t}^{\prime},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})\log(1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t))\right]$ (17)
$\stackrel{{\scriptstyle(i)}}{{=}}-\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})}{D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}+\frac{\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})}{1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}$
where step $(i)$ follows from standard differentiation of the logarithm function. We set (17) to be zero and get:
@@ E.1 Model Architecture
Our model architectures are based on the MDLM setting (Sahoo et al., 2024), which utilizes a Diffusion Transformer (DiT) (Peebles and Xie, 2023) with Rotary Position Embeddings (RoPE) (Su et al., 2024). We conduct experiments at two different scales. The first is a 169M parameter setting, and the second is a larger 424M parameter setting. The detailed hyperparameters for each are listed in Table 3 and Table 4, respectively.
@@ E.2 Dataset Details
We use the OpenWebText corpus for all experiments. The raw text is tokenized using the standard GPT-2 tokenizer. Following Sahoo et al. (2025), we concatenate all documents and pack them into fixed-length sequences of 1024 tokens, adding an <|endoftext|> token between documents. The final 100,000 documents of the corpus are reserved as the validation set.
For zero-shot evaluation, we assess the model’s generalization on seven diverse, out-of-domain datasets to measure its broader language understanding capabilities.
@@ E.3 Training and Distillation Hyperparameters
The detailed hyperparameters for teacher model pre-training and student model distillation are provided in Table 5. For our 169M models, we pre-trained the teacher model for 5 epochs, which took approximately 20 GPU hours on H100. The subsequent distillation process was completed in about one hour on a single H100 GPU. For the 424M models, the teacher was pre-trained for 7 epochs, requiring around 70 GPU hours on H100, while the corresponding distillation took approximately 2 hours on a single H100 GPU.
@@ E.4 Implementation Details
Our experiments were implemented using PyTorch. Key libraries and their versions are listed in Table 6. We utilized FlashAttention for optimized training performance. We plan to release our code and model checkpoints upon acceptance of the paper.
@@
FlashAttention 2.8.3
@@ E.5 Evaluation Details
Generative Perplexity (Gen PPL). We use the implementation from the transformers library to compute the perplexity of generated samples under a frozen GPT-2 Large model. Text is processed identically to the training data.
Entropy. To assess sample diversity, we generated 40 samples for each configuration. Each sample has a fixed length of 1024 tokens. We then computed the average token-level entropy across all generated samples.
Floating-Point Precision. All sampling and evaluation procedures were conducted using float64 precision. This is to avoid potential artifacts and misleading scores associated with lower-precision arithmetic, a practice highlighted as important for masked diffusion models (Sahoo et al., 2025).
@@ F.3 Implementation details of the ablation studies
The baseline model is a one-step generator, mapping a fully masked input $\mathbf{z}_{t}$ ($t=1$) directly to the final output $\mathbf{x}$. With this technique, we decompose the generation into a two-step process $\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}\rightarrow\mathbf{x}$ where $t=1$ and $\tau\in(0,1)$. The model first generates an intermediate state $\mathbf{z}_{\tau}$ at a randomly sampled time following $\pi(t)$ and then generates the final output $\mathbf{x}$ from $\mathbf{z}_{\tau}$. The ablation (‘w/o Score Decompose’) reverts to the direct, single-step generation.
Coupled Time $t$.
@@
This component synchronizes the time steps used for student generation and discriminator evaluation. In the standard setup, the intermediate denoising time $\tau$ for the student’s two-step generation ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$) and the corruption time $\tau^{\prime}$ for preparing the discriminator’s input ($\mathbf{x}\rightarrow\mathbf{z}_{\tau^{\prime}}$ for corrupted student generations, and $\mathbf{x}^{\prime}\rightarrow\mathbf{z}_{\tau^{\prime}}^{\prime}$ for corrupted teacher generations) are coupled, i.e., $\tau=\tau^{\prime}$. The ablation (‘w/o Coupled Time $t$’) decouples them by sampling $\tau$ and $\tau^{\prime}$ independently.
Weight Function $\omega(t)$ Correction.
@@
The gradient of the objective (5) is weighted by a time-dependent factor $\omega(t)$. The ablation (‘w/o $\omega(t)$ Correction’) uses a constant weight, i.e., $\omega(t)=1$. Our full model applies a theoretically-motivated correction based on the noise schedule’s signal rate $\alpha_{t}$, setting $\omega(t)\propto-\alpha^{\prime}_{t}/(1-\alpha_{t})$, which re-weights the objective based on the difficulty of the denoising step at time $t$.
Time Scheduler $\pi(\cdot)$ Weighting.
@@
This controls how $\tau$ is sampled and (optionally) whether we apply importance weighting, which also refers to the sampling distribution $\pi(\tau)$ for the intermediate time step $\tau$. The ablation setting (‘w/o $\pi(\tau)$ Weighting’) samples $\tau$ from a uniform distribution. Our method employs a non-uniform Beta distribution as $\pi(\tau)$ to focus the student’s training on specific, more informative intervals of the denoising trajectory. Specifically, we consider heavy-early ($\text{Beta}(2,5)$), and heavy-late ($\text{Beta}(5,2)$) schedules to improve sample quality. Practically, for smaller NFEs (e.g., 8 and 16 NFEs) we bias toward earlier times (e.g., $\text{Beta}(2,5)$); for larger NFEs (e.g., 32, 64, and 128 NFEs) we bias toward later times (e.g., $\text{Beta}(5,2)$).
Regularization.
@@
To stabilize training and maintain generalization, we introduce two regularization terms to the student’s loss. The first is a KL divergence loss, which aligns the student’s output logits with those of the frozen teacher at both steps of the generation process ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$ and $\mathbf{z}_{\tau}\rightarrow\mathbf{x}$, where $t=1$ and $\tau\in(0,1)$). Among Forward/Backward/Jeffreys implementations, we found forward KL consistently yielded the most stable optimization and best validation in our cases, so we adopt it for all ablations and apply it with a weight of $0.05$ in the student objective. The second term is an entropy bonus on the student’s output distribution to encourage exploration, which is weighted by $0.0005$. The ablation (‘w/o Regularization’) removes both the KL-divergence and entropy terms from the student’s objective.
Guided Inference.
@@
As elaborated in Section 4.4, this technique is applied only during sampling to leverage the trained discriminator and improve sample quality. We employ a hybrid strategy that divides the denoising process into two phases. For the first 50% of NFEs, we use gradient tilting ($M=1$), where the guidance scale $h$ is linearly increased from 0 to 30 to steer the generation. For the remaining 50% of NFEs, we switch to multi-candidate re-ranking ($h=0$), sampling $M=4$ candidates at each step and selecting the one with the highest reward. The ablation (‘w/o Guided Inference’) disables this process ($h=0,M=1$), reverting to standard unguided ancestral sampling.
F.4 Ablation study result analysis.
@@
The baseline model (without tricks) performs poorly, with a Gen PPL of 803.9 at 8 NFEs. As shown in Table 2, incorporating only the Score Decompose provides a moderate initial improvement. However, its indispensable role is starkly revealed in our leave-one-out study (Table 2). Removing this component from the full model results in a catastrophic performance degradation (PPL>30,000>30,000). This indicates a critical interplay: while a single-step gradient estimation is unstable within the complex training landscape created by our other optimizations, the two-step decomposition is essential for stabilizing the entire framework.
Time Coupling and Loss Weighting.
@@
The most significant performance gain is achieved by introducing Coupled Time $t$, which dramatically reduces the 8-step PPL from 667.8 to 101.0. This highlights the importance of aligning the temporal context between the reward signal and the score function. Loss reweighting then refines convergence: the $\omega(t)$ Correction improves mid/late-step budgets (e.g., $32{\to}64{\to}128$ NFEs: $48.4{\to}31.7{\to}25.3{\to}21.0$), while $\pi(t)$ Weighting is especially effective at $16$ NFEs ($75.6{\to}44.0$), with neutral to slight trade-offs elsewhere. The leave-one-out analysis confirms that removing any of these components leads to a severe degradation in performance.
The Dual Role of Regularization.
@@
Our study reveals that KL/entropy regularization plays a dual role depending on the sampling budget: it is helpful at very small budgets, where discretization error might be severe. For very small NFEs (e.g., 8 NFEs), where large discretization errors can destabilize training, regularization acts as a crucial stabilizer. The full model with regularization outperforms the version without it (PPL of 62.2 vs. 88.3). However, for more NFEs ($\geq 16$), Table 2 shows that removing regularization yields superior results, achieving the best PPLs in these settings (e.g., 30.99 vs. 38.19 at 16 NFEs). This suggests that for finer sampling schedules, the implicit stability of the chain is sufficient, and strong explicit regularization can over-constrain the model, leading to an accumulation of bias that harms the final generation quality.
Guided Inference.
@@
Guidance is most effective at small NFEs, where it markedly lowers PPL. Specifically, at 8 NFEs, guidance reduces PPL from 88.3 to 62.2, a relative improvement of $29.6\%$. This trend continues at 16 NFEs, where PPL drops by $13.2\%$ (from 44.0 to 38.2), and at 32 NFEs, where it decreases by $12.0\%$ (from 28.4 to 25.0). Conversely, at high NFEs, the effect on PPL is negligible (e.g., from $21.95$ to $21.91$ at 64 NFEs). In this regime, however, guidance substantially improves sample diversity. We observe that entropy increases from 5.06 to 5.15 at 64 NFEs and from 5.00 to 5.15 at 128 NFEs, indicating that guidance can enhance variety without sacrificing quality. This pattern matches our hybrid schedule: early gradient tilting improves accuracy at small budgets, while late multi-candidate re-ranking expands support and boosts diversity at larger budgets.

### v2

@@ Appendix C Student Objective
Our aim is to train a masked-diffusion student whose forward-time marginals match those of a teacher across the entire time horizon. We adopt the IKL between the student and teacher forward marginals $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ as the training objective. In the masked setting, the forward-time corruption kernel is a forward absorbing process, e.g.,
@@
$\nabla_{\nu}\mathcal{L}(\nu)\approx\frac{1}{N}\sum_{i=1}^{N}\left[\frac{\omega\left(t_{i}\right)}{\pi\left(t_{i}\right)}\cdot R\left(\mathbf{z}_{i},t_{i}\right)\cdot\nabla_{\nu}\log\mathbf{p}_{\nu}\left(\mathbf{z}_{t}=\mathbf{m},t=1\right)\right]$ (15)
@@ Appendix D Auxiliary Discriminator for Density Ratio Estimation
Our student objective, as derived in Appendix C, relies on the reward term $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$, which is intractable due to the unknown marginal distributions $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. To overcome this, we introduce a tractable estimator for the density ratio $\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t})}{\mathbf{q}_{\theta}(\mathbf{z}_{t})}$ by training an auxiliary discriminator network. This approach is inspired by the principles of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Wang et al., 2023) and has been successfully applied in Wang et al. (2025). We hereby adapt the following lemma to MDM setting.
For clarity, we adopt a time-agnostic notation where the corruption level is parameterized by the masking ratio $t\in[0,1]$, which corresponds to the schedule $\alpha_{t}$ via $t=1-\alpha_{t}$. We also denote $\mathbf{z}_{t}\in\mathcal{V}^{L}$ as a sequence with a mask ratio at time $t$.
Lemma D.1 (Density Ratio Representation, Restatement of Lemma 4.2).
Let $\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ and $\mathbf{q}_{\nu}(\mathbf{z}_{t},t)$ be the teacher and student model, respectively, over the discrete state space $\mathcal{V}^{L}$ at a mask ratio $\alpha_{t}$. Consider a discriminator $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$, which outputs a probability for each position $\ell\in\{1,\dots,L\}$. The discriminator $D_{\lambda}$ is trained to minimize the objective
@@
We here provide a detailed derivation for the optimal discriminator in the context of our discrete state space. The objective $\mathcal{L}_{D}(\lambda)$ is an expectation over the discrete random variable $\mathbf{z}_{t}$. We can express this expectation as a summation over all possible sequences $\mathbf{z}_{t}\in\mathcal{V}^{L}$:
$\mathcal{L}_{D}(\lambda)=\sum_{\mathbf{z}_{t}\in\mathcal{V}^{L}}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t},t)\log D_{\lambda}(\mathbf{z}_{t},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t},t)\log(1-D_{\lambda}(\mathbf{z}_{t},t))\right]$
@@
$\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}\mathcal{L}_{D}(\lambda)$ $=\frac{\partial}{\partial D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}\left[-\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})\log D_{\lambda}(\mathbf{z}_{t}^{\prime},t)-\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})\log(1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t))\right]$ (17)
$\stackrel{{\scriptstyle}}{{=}}-\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t}^{\prime},t^{\prime})}{D_{\lambda}(\mathbf{z}_{t}^{\prime},t)}+\frac{\mathbf{q}_{\theta}(\mathbf{z}_{t}^{\prime},t^{\prime})}{1-D_{\lambda}(\mathbf{z}_{t}^{\prime},t)},$
which follows from standard differentiation of the logarithm function. We set (17) to be zero and get:
@@ E.1 Model Architecture
Our model architectures are based on the MDLM setting (Sahoo et al., 2024), which utilizes a Diffusion Transformer (Peebles and Xie, 2023) with Rotary Position Embeddings (RoPE) (Su et al., 2024). We conduct experiments at two different scales. The first is a 169M parameter setting, and the second is a larger 424M parameter setting. The detailed hyperparameters for each are listed in Table 3 and Table 4, respectively.
@@ E.2 Dataset Details
We use the OpenWebText corpus for experiments. The raw text is tokenized using the standard GPT-2 tokenizer. Following Sahoo et al. (2025), we concatenate all documents and pack them into fixed-length sequences of 1024 tokens, adding an <|endoftext|> token between documents. The final 100,000 documents of the corpus are reserved as the validation set.
For zero-shot evaluation, we assess the model’s generalization on the following seven diverse, out-of-domain datasets to measure its broader language understanding capabilities.
@@ E.3 Training and Distillation Hyperparameters
The detailed hyperparameters for teacher model pre-training and student model distillation are provided in Table 5. For our 169M models, we pre-trained the teacher model for 5 epochs. The subsequent distillation process was completed in about one hour on a single H100 GPU. For the 424M models, the teacher was pre-trained for 7 epochs, while the corresponding distillation took approximately 2 hours on a single H100 GPU.
@@ E.4 Implementation Details
Our experiments were implemented using PyTorch. Key libraries and their versions are listed in Table 6. We plan to release our code and model checkpoints upon acceptance of the paper.
@@
Tokenizers 0.21.4
NumPy 2.3.2
SciPy 1.16.1
scikit-learn 1.7.1
Pandas 2.3.2
CUDA 12.1.0
cuDNN 9.1.0
Triton 3.1.0
@@ E.5 Evaluation Details
Generative Perplexity: We use the implementation from the transformers library to compute the perplexity of generated samples under a frozen GPT-2 Large model. Text is processed identically to the training data.
Entropy: To assess sample diversity, we generated 40 samples for each configuration. Each sample has a fixed length of 1024 tokens. We then computed the average token-level entropy across all generated samples.
Floating-Point Precision: All sampling and evaluation procedures were conducted using float64 precision. This is to avoid potential artifacts and misleading scores associated with lower-precision arithmetic, a practice highlighted as important for dLLMs (Sahoo et al., 2025).
@@ F.3 Implementation Details of the Ablation Studies
Score decompose. The baseline model is a one-step generator, mapping a fully masked input $\mathbf{z}_{t}$ ($t=1$) directly to the final output $\mathbf{x}$. With this technique, we decompose the generation into a two-step process $\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}\rightarrow\mathbf{x}$ where $t=1$ and $\tau\in(0,1)$. The model first generates an intermediate state $\mathbf{z}_{\tau}$ at a randomly sampled time following $\pi(t)$ and then generates the final output $\mathbf{x}$ from $\mathbf{z}_{\tau}$. The ablation (‘w/o Score Decompose’) reverts to the direct, single-step generation.
@@
Coupled time $t$. This component synchronizes the time steps used for student generation and discriminator evaluation. In the standard setup, the intermediate denoising time $\tau$ for the student’s two-step generation ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$) and the corruption time $\tau^{\prime}$ for preparing the discriminator’s input ($\mathbf{x}\rightarrow\mathbf{z}_{\tau^{\prime}}$ for corrupted student generations, and $\mathbf{x}^{\prime}\rightarrow\mathbf{z}_{\tau^{\prime}}^{\prime}$ for corrupted teacher generations) are coupled, i.e., $\tau=\tau^{\prime}$. The ablation (‘w/o Coupled Time $t$’) decouples them by sampling $\tau$ and $\tau^{\prime}$ independently.
@@
Weight function $\omega(t)$ correction. The gradient of the objective (5) is weighted by a time-dependent factor $\omega(t)$. The ablation (‘w/o $\omega(t)$ Correction’) uses a constant weight, i.e., $\omega(t)=1$. Our full model applies a theoretically-motivated correction based on the noise schedule’s signal rate $\alpha_{t}$, setting $\omega(t)\propto-\alpha^{\prime}_{t}/(1-\alpha_{t})$, which re-weights the objective based on the difficulty of the denoising step at time $t$.
@@
Time scheduler $\pi(\cdot)$ weighting. This controls how $\tau$ is sampled and (optionally) whether we apply importance weighting, which also refers to the sampling distribution $\pi(\tau)$ for the intermediate time step $\tau$. The ablation setting (‘w/o $\pi(\tau)$ Weighting’) samples $\tau$ from a uniform distribution. Our method employs a non-uniform Beta distribution as $\pi(\tau)$ to focus the student’s training on specific, more informative intervals of the denoising trajectory. Specifically, we consider heavy-early ($\text{Beta}(2,5)$), and heavy-late ($\text{Beta}(5,2)$) schedules to improve sample quality. Practically, for smaller NFEs (e.g., 8 and 16 NFEs) we bias toward earlier times (e.g., $\text{Beta}(2,5)$); for larger NFEs (e.g., 32, 64, and 128 NFEs) we bias toward later times (e.g., $\text{Beta}(5,2)$).
@@
Regularization. To stabilize training and maintain generalization, we introduce two regularization terms to the student’s loss. The first is a KL divergence loss, which aligns the student’s output logits with those of the frozen teacher at both steps of the generation process ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$ and $\mathbf{z}_{\tau}\rightarrow\mathbf{x}$, where $t=1$ and $\tau\in(0,1)$). Among Forward/Backward/Jeffreys implementations, we found forward KL consistently yielded the most stable optimization and best validation in our cases, so we adopt it for all ablations and apply it with a weight of $0.05$ in the student objective. The second term is an entropy bonus on the student’s output distribution to encourage exploration, which is weighted by $0.0005$. The ablation (‘w/o Regularization’) removes both the KL-divergence and entropy terms from the student’s objective.
@@
Guided inference. As elaborated in Section 4.4, this technique is applied only during sampling to leverage the trained discriminator and improve sample quality. We employ a hybrid strategy that divides the denoising process into two phases. For the first 50% of NFEs, we use gradient tilting ($M=1$), where the guidance scale $h$ is linearly increased from 0 to 30 to steer the generation. For the remaining 50% of NFEs, we switch to multi-candidate re-ranking ($h=0$), sampling $M=4$ candidates at each step and selecting the one with the highest reward. The ablation (‘w/o Guided Inference’) disables this process ($h=0,M=1$), reverting to standard unguided ancestral sampling.
F.4 Ablation Study Result Analysis
@@
Baseline and score decomposition. The baseline model (without tricks) performs poorly, with a Gen PPL of 803.9 at 8 NFEs. As shown in Table 2, incorporating only the Score Decompose provides a moderate initial improvement. However, its indispensable role is starkly revealed in our leave-one-out study (Table 2). Removing this component from the full model results in a catastrophic performance degradation (PPL>30,000>30,000). This indicates a critical interplay: while a single-step gradient estimation is unstable within the complex training landscape created by our other optimizations, the two-step decomposition is essential for stabilizing the entire framework.
@@
Time coupling and loss weighting. The most significant performance gain is achieved by introducing Coupled Time $t$, which dramatically reduces the 8-step PPL from 667.8 to 101.0. This highlights the importance of aligning the temporal context between the reward signal and the score function. Loss reweighting then refines convergence: the $\omega(t)$ Correction improves mid/late-step budgets (e.g., $32$ NFEs: from $48.4$ to $31.7$; $64$ NFEs: from $35.8$ to $25.3$), while $\pi(t)$ Weighting is especially effective at $16$ NFEs ($75.6{\to}44.0$), with neutral to slight trade-offs elsewhere. The leave-one-out analysis confirms that removing any of these components leads to a severe degradation in performance.
@@
The dual role of regularization. Our study reveals that KL/entropy regularization plays a dual role depending on the sampling budget: it is helpful at very small budgets, where discretization error might be severe. For very small NFEs (e.g., 8 NFEs), where large discretization errors can destabilize training, regularization acts as a crucial stabilizer. The full model with regularization outperforms the version without it (PPL of 62.2 vs. 88.3). However, for more NFEs ($\geq 16$), Table 2 shows that removing regularization yields superior results, achieving the best PPLs in these settings (e.g., 30.99 vs. 38.19 at 16 NFEs). This suggests that for finer sampling schedules, the implicit stability of the chain is sufficient, and strong explicit regularization can over-constrain the model, leading to an accumulation of bias that harms the final generation quality.
@@
Guided inference. Guidance is most effective at small NFEs, where it markedly lowers PPL. Specifically, at 8 NFEs, guidance reduces PPL from 88.3 to 62.2, a relative improvement of $29.6\%$. This trend continues at 16 NFEs, where PPL drops by $13.2\%$ (from 44.0 to 38.2), and at 32 NFEs, where it decreases by $12.0\%$ (from 28.4 to 25.0). Conversely, at high NFEs, the effect on PPL is negligible (e.g., from $21.95$ to $21.91$ at 64 NFEs). In this regime, however, guidance substantially improves sample diversity. We observe that entropy increases from 5.06 to 5.15 at 64 NFEs and from 5.00 to 5.15 at 128 NFEs, indicating that guidance can enhance variety without sacrificing quality. This pattern matches our hybrid schedule: early gradient tilting improves accuracy at small budgets, while late multi-candidate re-ranking expands support and boosts diversity at larger budgets.